# mode = titleCMD-HAR

### v3

@@ 1 Introduction
Human Activity Recognition (HAR) aims to identify the types of activities performed by humans based on information received from different devices, such as cameras and motion sensors [1]. HAR has been widely used in various real-world applications, especially in healthcare [2] and exercise assessment [3]. Based on data type, HAR systems can be split into two major categories: video-based HAR and sensor-based HAR. However, video-based systems recognize behaviors through captured images or videos, facing societal and technical challenges, such as privacy concerns, dependence on environmental lighting, resolution constraints, and the high cost and complexity associated with video processing algorithms. In contrast, sensor-based HAR has been widely used because of their ease of installation, and non-intrusiveness [4]. Until now, mobile devices have brought about a new way of life and have become an indispensable part of people’s daily lives [5].
@@ 2.1 Feature Extraction of Complex Activities
Prior to the advent of deep learning, HAR systems were predominantly based on manual feature engineering. The spatiotemporal interest point (STIP) method proposed by Laptev [6] in 2005 exemplifies this approach, utilizing spatiotemporal and geometric information for feature representation. However, such methods required extensive domain expertise and demonstrated limited efficacy in complex activity recognition. Subsequent advances introduced data-driven approaches, notably Simonyan and Zisserman’s two-stream CNN (2D-CNN) [7], which fused spatial features from single frames with temporal features from optical flow maps. Although effective for temporal modeling, 2D-CNNs exhibited limitations in capturing long-range dependencies. This led to innovations such as P-CNN [8], which incorporated human body part trajectories to better model complex motions.
@@ 2.2 Disentangled Representation Learning
Disentangled representation learning seeks to isolate key explanatory factors in data generation while preserving the interpretability of the model, a critical requirement for HAR systems. The theoretical foundation was established by Bengio et al. [9], who posited that disentangled representations enable better generalization through separation of latent factors. Practical implementations emerged with Higgins’ $\beta$-VAE [10], which introduced a tunable hyperparameter to balance reconstruction accuracy and factor disentanglement. Kim and Mnih’s FactorVAE [11] further improved disentanglement efficacy through dedicated regularization terms.
@@ 2.3 Modal Heterogeneity
Current approaches to modal heterogeneity face three principal challenges: 1) Data imputation techniques such as GANs, which have been employed to address issues like inconsistent missing data patterns in HAR, rely on high-quality training data and exhibit limited generalizability [14]. 2) Conventional grouping strategies restrict cross-modal knowledge transfer [15]. 3) Existing knowledge-sharing paradigms fail to fully integrate multi-scale features [16]. Gradient modulation methods provide a promising solution to address modal heterogeneity in HAR, significantly enhancing multimodal information fusion. Peng et al. [17] proposed the dynamic gradient modulation method, which adjusts modal contributions during optimization, though restricted to late fusion architectures. Li et al. [18] explored an adaptive gradient modulation (AGM) method based on Shapley values, which is applicable to complex fusion strategies and introduces a novel modality competition intensity metric to reveal the underlying mechanisms of modality competition. Guo et al. [19] introduced the classifier-guided gradient modulation (CGGM) method, which evaluates modality utility using lightweight modal classifiers, integrating both gradient magnitude and direction. Compared to existing methods, CGGM is not restricted by loss functions, optimizers, or the number of modalities, but the introduction of additional classifiers increases computational overhead.
@@ 3.3 Spatiotemporal Disentanglement Mechanism
$L=\lambda_{1}{||\mathbf{F}_{S}-\mathbf{F}_{T}||_{2}}+\lambda_{2}{||\mathbf{F}_{P}-\mathbf{F}_{S}||_{2}}+\lambda_{3}{||\mathbf{F}_{P}-\mathbf{F}_{T}||_{2}},$ (2)
@@ 4.1 Dataset used in the experiments
WISDM [20]: Smartphone accelerometers recorded data at 20 Hz from 36 participants, capturing 18 activities such as walking, running, sitting, and climbing stairs. UCI-HAR [21]: This dataset was collected using waist-mounted smartphones with accelerometers and gyroscopes at 50 Hz, involving 30 participants and daily activities such as walking, running and sitting. PAMAP2 [22]: With wearable sensors collecting data at 100 Hz, 9 participants performed 18 activities, including walking, running, and cycling. USC-HAD [23]: Collected using wearable sensors at 50 Hz, this dataset includes 12 activities performed by 14 participants in their daily routines. OPPORTUNITY [24]: This dataset features 20 activities recorded with wearable and environmental sensors at multiple frequencies, designed for multimodal HAR research. UniMiB-SHAR [25]: Collected using smartphone accelerometers at 50 Hz, with 50 participants performing 30 activities (including daily behaviors and fall scenarios).
@@ 5.1 Ablation Experiment
CNN[26] 85.09 0.85 0.79 75.92 0.76 0.65 96.50 0.94 0.94 93.18 0.93 0.93
LSTM[27] 85.91 0.86 0.79 84.25 0.84 0.76 99.60 0.99 0.99 88.12 0.88 0.88
CNN-GRU[28] 74.86 0.73 0.72 83.07 0.83 0.74 81.11 0.81 0.80 90.80 0.91 0.91
LSTM-CNN[27] 81.08 0.81 0.75 87.84 0.88 0.81 98.97 0.98 0.98 89.04 0.89 0.89
Rev-Attention[29] 59.71 0.59 0.44 75.22 0.75 0.62 98.98 0.98 0.98 91.92 0.92 0.92
SE-Res2Net[30] 82.96 0.83 0.75 86.12 0.79 0.86 88.42 0.87 0.87 90.96 0.91 0.91
ResNext[31] 83.05 0.83 0.75 87.57 0.87 0.81 91.68 0.90 0.90 89.76 0.89 0.89
Gated-Res2Net[32] 81.98 0.82 0.75 86.86 0.87 0.80 89.62 0.88 0.88 86.90 0.85 0.85
MAG-Res2Net[33] 87.39 0.88 0.80 86.40 0.86 0.79 89.89 0.89 0.89 91.56 0.92 0.91
HAR-CE[34] 83.05 0.83 0.74 86.43 0.86 0.79 91.89 0.91 0.91 91.35 0.91 0.91
ELK[35] 82.39 0.83 0.73 88.19 0.88 0.83 93.12 0.93 0.92 91.50 0.91 0.91
DanHAR[36] 86.16 0.86 0.77 86.00 0.86 0.78 94.35 0.94 0.93 91.44 0.91 0.92
RepMobile [37] 72.47 0.72 0.69 83.66 0.83 0.74 90.30 0.91 0.90 92.68 0.93 0.91
SACNN-GA [38] 75.53 0.74 0.68 84.66 0.84 0.69 91.75 0.92 0.91 87.85 0.89 0.88
@@ 5.3 Actual Deployment
To evaluate efficiency of the model, we conducted inference delay tests, measuring the time taken to infer a single sample and 100 consecutive samples. We followed Zhang’s inference delay testing method, using a sliding window approach to evaluate sensor data [39], selecting 10-second action clips with a sliding window step of 95 % of the window length (500 ms). Due to improved device performance, most models easily met the required standard. Considering that the WISDM dataset typically uses 4-second segments with a 200-millisecond step, we applied this method to each data set to ensure the model processes the next segment in time.

### v4

@@ 1 Introduction
Human Activity Recognition (HAR) aims to identify the types of activities performed by humans based on information received from different devices, such as cameras and motion sensors [yang2022cross]. HAR has been widely used in various real-world applications, especially in healthcare [zhou2020deep] and exercise assessment [zhou2023swarm]. Based on data type, HAR systems can be split into two major categories: video-based HAR and sensor-based HAR. However, video-based systems recognize behaviors through captured images or videos, facing societal and technical challenges, such as privacy concerns, dependence on environmental lighting, resolution constraints, and the high cost and complexity associated with video processing algorithms. In contrast, sensor-based HAR has been widely used because of their ease of installation, and non-intrusiveness [essa2023temporal]. Until now, mobile devices have brought about a new way of life and have become an indispensable part of people’s daily lives [yin2024systematic].
@@ 2.1 Feature Extraction of Complex Activities
Prior to the advent of deep learning, HAR systems were predominantly based on manual feature engineering. The spatiotemporal interest point (STIP) method proposed by Laptev [1238378] in 2005 exemplifies this approach, utilizing spatiotemporal and geometric information for feature representation. However, such methods required extensive domain expertise and demonstrated limited efficacy in complex activity recognition. Subsequent advances introduced data-driven approaches, notably Simonyan and Zisserman’s two-stream CNN (2D-CNN) [DBLP:journals/corr/SimonyanZ14], which fused spatial features from single frames with temporal features from optical flow maps. Although effective for temporal modeling, 2D-CNNs exhibited limitations in capturing long-range dependencies. This led to innovations such as P-CNN [7410725], which incorporated human body part trajectories to better model complex motions.
@@ 2.2 Disentangled Representation Learning
Disentangled representation learning seeks to isolate key explanatory factors in data generation while preserving the interpretability of the model, a critical requirement for HAR systems. The theoretical foundation was established by Bengio et al. [10.1109/TPAMI.2013.50], who posited that disentangled representations enable better generalization through separation of latent factors. Practical implementations emerged with Higgins’ $\beta$-VAE [Higgins2016betaVAELB], which introduced a tunable hyperparameter to balance reconstruction accuracy and factor disentanglement. Kim and Mnih’s FactorVAE [pmlr-v80-kim18b] further improved disentanglement efficacy through dedicated regularization terms.
@@ 2.3 Modal Heterogeneity
Current approaches to modal heterogeneity face three principal challenges: 1) Data imputation techniques such as GANs, which have been employed to address issues like inconsistent missing data patterns in HAR, rely on high-quality training data and exhibit limited generalizability [hussein2022robust]. 2) Conventional grouping strategies restrict cross-modal knowledge transfer [palechor2022semi]. 3) Existing knowledge-sharing paradigms fail to fully integrate multi-scale features [guo2025multi]. Gradient modulation methods provide a promising solution to address modal heterogeneity in HAR, significantly enhancing multimodal information fusion. Peng et al. [Peng2022BalancedML] proposed the dynamic gradient modulation method, which adjusts modal contributions during optimization, though restricted to late fusion architectures. Li et al. [li2023boosting] explored an adaptive gradient modulation (AGM) method based on Shapley values, which is applicable to complex fusion strategies and introduces a novel modality competition intensity metric to reveal the underlying mechanisms of modality competition. Guo et al. [guo2024classifier] introduced the classifier-guided gradient modulation (CGGM) method, which evaluates modality utility using lightweight modal classifiers, integrating both gradient magnitude and direction. Compared to existing methods, CGGM is not restricted by loss functions, optimizers, or the number of modalities, but the introduction of additional classifiers increases computational overhead.
@@ 3.3 Spatiotemporal Disentanglement Mechanism
[…]
@@ 4.1 Dataset used in the experiments
WISDM [kwapisz2011activity]: Smartphone accelerometers recorded data at 20 Hz from 36 participants, capturing 18 activities such as walking, running, sitting, and climbing stairs. UCI-HAR [anguita2013public]: This dataset was collected using waist-mounted smartphones with accelerometers and gyroscopes at 50 Hz, involving 30 participants and daily activities such as walking, running and sitting. PAMAP2 [reiss2012introducing]: With wearable sensors collecting data at 100 Hz, 9 participants performed 18 activities, including walking, running, and cycling. USC-HAD [zhang2012usc]: Collected using wearable sensors at 50 Hz, this dataset includes 12 activities performed by 14 participants in their daily routines. OPPORTUNITY [chavarriaga2013opportunity]: This dataset features 20 activities recorded with wearable and environmental sensors at multiple frequencies, designed for multimodal HAR research. UniMiB-SHAR [micucci2017unimib]: Collected using smartphone accelerometers at 50 Hz, with 50 participants performing 30 activities (including daily behaviors and fall scenarios).
@@ 5.1 Ablation Experiment
CNN[zeng2014convolutional] 85.09 0.85 0.79 75.92 0.76 0.65 96.50 0.94 0.94 93.18 0.93 0.93
LSTM[xia2020lstm] 85.91 0.86 0.79 84.25 0.84 0.76 99.60 0.99 0.99 88.12 0.88 0.88
CNN-GRU[dua2021multi] 74.86 0.73 0.72 83.07 0.83 0.74 81.11 0.81 0.80 90.80 0.91 0.91
LSTM-CNN[xia2020lstm] 81.08 0.81 0.75 87.84 0.88 0.81 98.97 0.98 0.98 89.04 0.89 0.89
Rev-Attention[pramanik2023transformer] 59.71 0.59 0.44 75.22 0.75 0.62 98.98 0.98 0.98 91.92 0.92 0.92
SE-Res2Net[gao2019res2net] 82.96 0.83 0.75 86.12 0.79 0.86 88.42 0.87 0.87 90.96 0.91 0.91
ResNext[mekruksavanich2022deep] 83.05 0.83 0.75 87.57 0.87 0.81 91.68 0.90 0.90 89.76 0.89 0.89
Gated-Res2Net[yang2020gated] 81.98 0.82 0.75 86.86 0.87 0.80 89.62 0.88 0.88 86.90 0.85 0.85
MAG-Res2Net[liu2023mag] 87.39 0.88 0.80 86.40 0.86 0.79 89.89 0.89 0.89 91.56 0.92 0.91
HAR-CE[CE] 83.05 0.83 0.74 86.43 0.86 0.79 91.89 0.91 0.91 91.35 0.91 0.91
ELK[ELK] 82.39 0.83 0.73 88.19 0.88 0.83 93.12 0.93 0.92 91.50 0.91 0.91
DanHAR[DanHAR] 86.16 0.86 0.77 86.00 0.86 0.78 94.35 0.94 0.93 91.44 0.91 0.92
RepMobile [yu2024repmobile] 72.47 0.72 0.69 83.66 0.83 0.74 90.30 0.91 0.90 92.68 0.93 0.91
SACNN-GA [sarkar2023human] 75.53 0.74 0.68 84.66 0.84 0.69 91.75 0.92 0.91 87.85 0.89 0.88
@@ 5.3 Actual Deployment
To evaluate efficiency of the model, we conducted inference delay tests, measuring the time taken to infer a single sample and 100 consecutive samples. We followed Zhang’s inference delay testing method, using a sliding window approach to evaluate sensor data [NNU2023], selecting 10-second action clips with a sliding window step of 95 % of the window length (500 ms). Due to improved device performance, most models easily met the required standard. Considering that the WISDM dataset typically uses 4-second segments with a 200-millisecond step, we applied this method to each data set to ensure the model processes the next segment in time.